# SlideAgent: Hierarchical Agentic Framework for Multi-Page Visual Document Understanding

### v2

@@ 3.2 Performance with Ground-truth Pages
When ground-truth pages are provided (Table 11/12), the performance gap narrows. As all models receive the exact pages with the answers, noise introduced in retrieval is eliminated. Under this oracle setting, SlideAgent still improves over the base model (+7.7 overall and +12.5 numeric on SlideVQA), demonstrating the effectiveness of element-level retrieval.
Text-based Retrievers MRR Recall@1 nDCG@1 Recall@3 Hit@3 nDCG@3
BM25 Robertson et al. (2004) 59.0 51.5 52.0 63.0 65.3 57.7
w/ SA 63.9 +4.9 54.9 +3.4 56.6 +4.6 67.1 +4.1 68.8 +3.5 62.6 +4.9
BGE Xiao et al. (2023) 70.1 56.1 60.9 75.8 78.1 69.2
w/ SA 72.3 +2.2 58.4 +2.3 63.2 +2.3 77.4 +1.6 80.3 +2.2 71.3 +2.1
SFR Meng et al. (2024) 70.1 56.1 60.9 75.8 78.1 69.2
w/ SA 76.5 +6.4 59.9 +3.8 69.4 +8.5 77.3 +1.5 81.8 +3.7 73.2 +4.0
Multimodal Retrievers MRR Recall@1 NDCG@1 Recall@3 Hit@3 NDCG@3
SigLIP2 Tschannen et al. (2025) 26.7 15.9 18.0 31.0 34.0 25.0
w/ SA 28.0 +1.3 16.3 +0.4 18.0 +0.0 32.3 +1.3 35.5 +1.5 26.1 +1.1
COLPALI Faysse et al. (2024) 82.1 68.2 75.5 78.9 84.0 77.4
w/ SA 82.9 +0.8 70.4 +2.2 76.2 +0.7 88.6 +9.7 90.1 +6.1 83.1 +5.7
VisRAG Yu et al. (2025) 76.0 63.3 68.6 82.3 84.1 76.0
w/ SA 79.7 +3.7 66.3 +3.0 71.6 +3.0 85.5 +3.2 87.7 +3.6 79.4 +3.4
Table 4: Artifacts generated by SlideAgent, particularly $\hat{Q}$ and $\mathcal{M}_{p}$, improves page-level retrieval performance, especially for text-based retrievers.
@@ 3.3 Effectiveness of Knowledge Construction
We evaluate whether hierarchical knowledge representations ($\mathcal{K}$) improve retrieval beyond end-to-end QA. Specifically, we test page-level retrieval using generated subqueries $\hat{Q}$ and page knowledge $\mathcal{K}_{p}$ from SlideAgent. Table 4 shows consistent gains across both text-based retrievers (BM25 Robertson et al. (2004), BGE Xiao et al. (2023), SFR Meng et al. (2024)) and multimodal retrievers (COLPALI Faysse et al. (2024), VisRAG Yu et al. (2025), SigLIP2 Tschannen et al. (2025)).
@@
COLPALI Faysse et al. (2024) improves slightly in ranking quality (+0.8 MRR) but substantially in coverage (+9.7 Recall@3), indicating that structured subqueries help it surface more relevant pages even if they are not ranked at the very top. VisRAG Yu et al. (2025) achieves a larger ranking boost (+3.7 MRR), suggesting that multimodal LLMs still benefit from textual guidance in aligning queries to page content. SigLIP2 shows minimal gains (+1.3 MRR), likely because it is optimized for natural image domains and transfers less effectively to document-style inputs.
Figure 5: Example answers generated by SlideAgent. Agents at different levels work together to provide comprehensive responses.
@@ 3.4 Ablation Studies
We analyze each design choice by removing global agent (w/o G), page agent (w/o P), element agent (w/o E), and subquery generation (w/o S). Figures 3/6 summarize results across proprietary (GPT-4o) and open-source (InternVL3-8B) backbones.
@@
Removing the page agent causes the steepest degradation–GPT-4o drops –6.3 overall (–9.5 numeric) and InternVL3-8B –8.8 overall. The page agent integrates global themes $\mathcal{K}_{g}$ and sequential context $\mathcal{K}_{p}^{i-1}$ (Eq. 1), enabling cross-slide coherence and multi-hop reasoning. Its absence breaks this structural bridge between global context and page details. Results for SlideAgent w/o E, w/o G, and w/o S are in Appendix D.1.
@@ 4 Related Work
In this section, we discuss relevant works in visual document understanding and Multimodal LLMs. A more detailed description is in Appendix B.
@@
5 Conclusion
We present SlideAgent, a hierarchical agentic framework that leverages specialized agents at multiple levels for fine-grained infographics understanding, particularly for slide decks. SlideAgent demonstrates significant performance gains and strong generalizability across both proprietary and open-source models.
@@ 6 Limitations
We did not evaluate element-level retrieval due to the lack of corresponding annotations in existing datasets and the difficulty of defining precise boundaries among elements. Benchmarks in visual document understanding could include such annotations to enable finer-grained analysis.
Retrieval Method.
Our framework supports both textual and multimodal retrievers. We primarily adopt text-based retrieval for efficiency, though future work could explore multimodal retrievers or domain-specific strategies tailored to slide decks.
7 Ethical Considerations
Data Privacy, Consent, and Intellectual Property.
Content Reliability and User Responsibility.
@@ 7 Ethical Considerations
Appendix A Details Method
A.1 Element Merging
@@
Notation Description
$p_{i},v_{i}$ A slide page and its visual content
$\mathcal{P}=\{p_{1},p_{2},\dots\}$ Set of pages in the slide deck
$q$ User query
$\hat{Q}$ Set of subqueries generated from the original query $q$
$a$ The final answer generated by SlideAgent
$b_{j}$ Bounding box coordinates of an element within a slide
$t_{j}$ Type of the element (e.g., text, chart, image, table)
[…]
Table 6: Mathematical notation used throughout the paper.
Case Characteristics Example Agents
1-Global Understanding Asks about the overall theme, purpose, or general summary of the entire presentation. "What is the main topic of the presentation?", "What is this deck about?" Global
2-Fact-based Direct Query Asks for specific facts, data, or information from particular slides. "What is the revenue reported on slide 7?", "Which slide shows the product roadmap?" Page, Element
3-Multi-hop Reasoning Requires comparing information across multiple slides or elements. "Compare revenues in slide 5 and slide 10", "How do Sweden and Denmark compare?" Global, Page, Element
4-Layout / Visual Relationship Asks about visual relationships, positioning, or layout elements. "What does the diagram below the table on slide 12 illustrate?", "Is the color red used to denote negative performance?" Element
5-Uncertain If the query is unclear, use all agents to answer the query. \ Global, Page, Element
Appendix B Extended Related Work
B.1 Document and Infographics Understanding
B.1.1 Slide Deck Understanding
B.1.2 Slide Visual Question Answering
B.1.3 General Purpose LLMs
General-purpose large language models (MLLMs) such as GPT-4 Achiam et al. (2023), GPT-4o Hurst et al. (2024), Gemini Anil et al. (2023), LLaVA Liu et al. (2023), InternVL3 Zhu et al. (2025), and Visual CoT Shao et al. (2024) have significantly advanced visual reasoning across a wide range of tasks. Visual prompting methods like Set-of-Marks Yang et al. (2023) further enhance reasoning by augmenting inputs with structured visual annotations. While these open-source and closed-source MLLMs form the foundation for many comparisons, they often struggle when applied to slides, where precise grounding and fine-grained interpretation are required to parse heterogeneous elements such as charts, tables, and icons.
@@ B.1.3 General Purpose LLMs
This approach effectively reduces element fragmentation while preserving spatial relationships, enabling more accurate element-level descriptions for our hierarchical framework. Algorithm 1 provides the complete implementation details.
@@ B.2 Answer Matching
To determine answer equivalence, i.e. if two agents agree upon their opinions, we do fuzzy string matching between each pair of answers $a_{1},a_{2}$ using Normalized Levenshtein Similarity (NLS):
@@
where $\text{editdistance}(a_{1},a_{2})$ computes the Levenshtein distance (Lcvenshtcin, 1966) between the answers after tokenization. Two answers are considered equivalent if $\text{NLS}(a_{1},a_{2})\geq 0.75$.
Appendix C Experimental Setup
@@ C.1 Dataset Descriptions
  • SlideVQA (Tanaka et al., 2023) is a multi-modal, multi-image VQA dataset consisting of $\geq$ 2600 slide decks that require single-hop, multi-hop, and numerical reasoning. Each slide deck corresponds to multiple questions. The license is availble at  [^1] https://github.com/nttmdlab-nlp/SlideVQA/blob/main/LICENSE.
  • InfoVQA (Mathew et al., 2022) is a dataset that feature questions that require reasoning and arithmetic skills. It contains over 30,000 questions with over 5,400 images. The dataset is released under CC-BY license.
@@ C.3 Implementation Details
We used EasyOCR [^2] https://github.com/JaidedAI/EasyOCR and Docling (Team, 2024) for detecting textual and visual elements, respectively. Whenever applicable, the answers are generated using a temperature of 0.0 to ensure deterministic results.
@@
Appendix D Additional Experiments
D.1 Ablation Studies (Cont’d)
Element-level Reasoning Ensures Precision
Impact of Global Thematic Guidance
Removing the global agent yields the smallest drop (–2.8 GPT-4o; –3.7 InternVL3-8B), as lower-level agents already embed partial global context via $\mathcal{K}_{p}$ and $\mathcal{K}_{e}$ in knowledge construction (Eq. 1/2). However, responses become less aware of overarching themes.
Subquery Generation Strengthens Retrieval.
Appendix E Future Work
Element Parsing.
SlideAgent relies on advanced text and layout detection tools, which may struggle with visually complex or low-contrast designs (e.g., white charts on white backgrounds). Enhancing robustness in parsing could further improve overall reliability.
Modeling Element Relations.
SlideAgent currently treats elements independently. Extending it to explicitly model inter-element relations–for instance, via graph structures linking textual and visual components–offers a promising direction for capturing richer semantics, though at higher computational cost.
User-Centric Interaction.
While SlideAgent primarily enhances reasoning and retrieval for slide understanding, future work could make such agentic frameworks more user-centric through seamless integration into real-world user behaviors Wang et al. (2025b). Lightweight feedback mechanisms (e.g., contextual pop-ups or in-situ clarifications) could further improve interpretability and interaction without disrupting natural workflows or requiring excessive user input.
### Cumulative Summary of the Preceding Slides {Cumulative Summary}
Element Type <<text | image | chart | table | icon | button | etc.>>
Position on Slide <<e.g., top-right, centered, below title>>
Semantic Role <<What is the element trying to do or communicate?>>
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Raw Models
Gemini 2.0 86.3 81.0 90.3 59.7 60.0 59.6 78.1 77.8 88.9
Gemini 2.5 89.0 85.7 93.2 61.5 65.0 59.5 76.6 76.4 83.3
Gemini 2.5-lite 81.8 75.5 90.1 56.3 55.0 57.0 78.1 78.4 66.7
Claude 4.1 85.7 82.4 89.7 58.0 77.5 48.3 52.6 52.0 60.8
Claude 3.5 58.2 64.0 50.9 55.8 83.7 42.5 47.8 48.0 40.3
GPT-4o 79.4 71.9 86.4 64.5 77.1 58.0 83.0 84.3 80.1
Multimodal RAG and Agentic Methods
ViDoRAG 81.8 73.8 87.0 65.8 78.0 58.7 84.2 84.7 81.5
SlideAgent 87.1 84.4 90.6 68.7 82.5 61.5 85.8 85.9 85.6
Impr. +7.7 +12.5 +4.2 +4.1 +5.4 +3.5 +2.8 +1.5 +5.5
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Raw Models
Llama 3.2 11B 44.6 52.1 34.6 47.1 62.8 39.3 39.1 39.2 33.5
Phi3 78.3 69.1 91.6 53.9 67.4 47.2 63.8 63.7 65.1
Qwen2.5 7B 85.1 77.7 95.3 57.7 69.8 51.8 52.7 52.0 77.8
Qwen2.5 32B 87.4 82.6 93.7 49.6 62.8 43.2 69.5 69.6 65.1
LLaVA 1.5 7B 42.9 27.9 79.9 24.6 15.0 39.4 14.2 14.4 20.9
LLaVA 1.5 13B 46.7 29.5 83.1 29.2 17.5 46.6 23.8 20.3 41.2
LLaVA 1.6 7B 59.0 45.8 84.2 36.2 40.0 34.0 16.0 15.2 21.3
LLaVA 1.6 13B 62.3 48.2 87.1 54.7 62.5 49.5 35.2 30.7 67.6
InternVL3 8B 73.3 65.4 85.7 58.4 72.1 51.5 56.3 55.9 65.6
Baseline methods based on InternVL3-8B
ViDoRAG 76.3 68.1 89.9 61.3 75.1 53.9 58.4 58.7 67.3
SlideAgent 82.8 75.3 93.3 64.6 79.5 58.0 62.8 62.6 68.1
Impr. +9.5 +9.8 +7.6 +6.2 +7.4 +6.4 +6.5 +6.7 +2.5

### v3

@@ 3.2 Performance with Ground-truth Pages
When ground-truth pages are provided (Table 11/12), the performance gap narrows. As all models receive the exact pages with the answers, noise introduced in retrieval is eliminated. Under this oracle setting, SlideAgent still improves over the base model (+7.7+7.7% overall and +12.5+12.5% numeric on SlideVQA), demonstrating the effectiveness of element-level retrieval.
3.3 Error Analysis
We annotate 40 SlideVQA failure cases to quantify parsing-related error propagation. As shown in Table 5, only OCR Error and Tiny Visuals are directly attributable to parsing failures, accounting for 5/40 cases (12.5%). The dominant categories are Ambiguous Question (22.5%), Answer Mislocation (17.5%), and Valid Alternative Answer (15.0%), indicating that ambiguity, retrieval drift, and annotation subjectivity account for a larger share of failures than OCR/layout parsing.
Figure 5: Error analysis of 40 failure cases on SlideVQA. Only OCR Error and Tiny Visuals (12.5% combined) are directly attributable to parsing failures.
@@ 3.4 Effectiveness of Knowledge Construction
We evaluate whether hierarchical knowledge representations ($\mathcal{K}$) improve retrieval beyond end-to-end QA. Specifically, we test page-level retrieval using generated subqueries $\hat{Q}$ and page knowledge $\mathcal{K}_{p}$ from SlideAgent. Figure 6 shows significant gains across both text-based retrievers (BM25 Robertson et al. (2004), BGE Xiao et al. (2023), SFR Meng et al. (2024)) and multimodal retrievers (COLPALI Faysse et al. (2024), VisRAG Yu et al. (2025), SigLIP2 Tschannen et al. (2025)).
Figure 6: Artifacts generated by SlideAgent, particularly $\hat{Q}$ and $\mathcal{M}_{p}$, improves page-level retrieval performance, especially for text-based retrievers.
@@
COLPALI Faysse et al. (2024) improves slightly in ranking quality (+0.8 MRR) but substantially in coverage (+9.7 Recall@3), indicating that structured subqueries help it surface more relevant pages even if they are not ranked at the very top. VisRAG Yu et al. (2025) achieves a larger ranking boost (+3.7+3.7% MRR), suggesting that multimodal LLMs still benefit from textual guidance in aligning queries to page content. SigLIP2 shows minimal gains (+1.3+1.3% MRR), likely because it is optimized for natural image domains and transfers less effectively to document-style inputs.
@@ 3.5 Ablation Studies
We analyze each design choice by removing global agent (w/o G), page agent (w/o P), element agent (w/o E), and subquery generation (w/o S). Figures 7 summarize results across proprietary (GPT-4o) and open-source (InternVL3-8B) backbones.
@@
Removing the page agent causes the steepest degradation–GPT-4o drops –6.3 overall (–9.5 numeric) and InternVL3-8B –8.8 overall. The page agent integrates global themes $\mathcal{K}_{g}$ and sequential context $\mathcal{K}_{p}^{i-1}$ (Eq. 1), enabling cross-slide coherence and multi-hop reasoning. Its absence breaks this structural bridge between global context and page details.
Element-level Reasoning Ensures Precision
Impact of Global Thematic Guidance
Removing the global agent yields the smallest drop (-2.82.8% GPT-4o; –$3.7$ InternVL3-8B), as lower-level agents already embed partial global context via $\mathcal{K}_{p}$ and $\mathcal{K}_{e}$ in knowledge construction (Eq. 1/3). However, responses become less aware of overarching themes.
Subquery Generation Strengthens Retrieval.
@@ 3.6 Qualitative Answer Analysis
Figure 8: Example answers by SlideAgent. Agents at different levels collaborate for comprehensive responses.
@@ 4 Related Work
In this section, we discuss relevant works in visual document understanding and Multimodal LLMs. A more detailed description is in Appendix A.
@@
5 Conclusion & Future Work
We present SlideAgent, a hierarchical agentic framework that leverages specialized agents at multiple granularity levels for multi-page visual document understanding, particularly for slide decks. Future work can improve parsing robustness on low-contrast layouts, model inter-element relations via graphs, replace the first-$N$-pages heuristic with content-aware page selection for $\mathcal{K}_{g}$, and extend SlideAgent to multi-turn scenarios.
@@ 6 Limitations
1) Element-level Retrieval. We assess element-level grounding on a manually annotated subset of SlideVQA questions, providing direct evidence of fine-grained retrieval. However, element boundaries may vary across OCR and layout-parsing tools. Broader benchmarks with element-level annotations would enable more comprehensive evaluation. 2) Retrieval Method. Our framework supports both textual and multimodal retrieval. For efficiency, we primarily adopt text-based retrieval. Future work could explore multimodal retrievers or domain-specific retrieval strategies.
@@
Acknowledgements
The authors would like to thank Nishan Srishankar, Kelly Patel, and Manuela Veloso at J.P. Morgan AI Research for their valuable discussions and feedback on this work.
Appendix A Extended Related Work
A.1 Document and Infographics Understanding
A.1.1 Slide Deck Understanding
A.1.2 Slide Visual Question Answering
A.1.3 General Purpose LLMs
General-purpose large language models (MLLMs) such as GPT-4 Achiam et al. (2023), GPT-4o Hurst et al. (2024), Gemini Anil et al. (2023), LLaVA Liu et al. (2023), InternVL3 Zhu et al. (2025), and Visual CoT Shao et al. (2024) have significantly advanced visual reasoning across a wide range of tasks. Visual prompting methods like Set-of-Marks Yang et al. (2023) enhance reasoning by augmenting inputs with structured visual annotations. Multi-agent systems Wang et al. (2025b, 2026) have also enhanced the multi-perspective reasoning capabilities of LLM-powered systems. While these open-source and closed-source MLLMs form the foundation for many comparisons, they often struggle when applied to slides, where precise grounding and fine-grained interpretation are required to parse heterogeneous elements such as charts, tables, and icons.
Appendix B Additional Method Details
B.1 Element Merging
@@
This approach effectively reduces element fragmentation while preserving spatial relationships, enabling more accurate element-level descriptions for our hierarchical framework. Algorithm 1 (at the end of the appendix) provides the complete implementation details.
@@ B.2 Answer Matching
To determine answer equivalence, i.e., if two agents agree upon their opinions, we do fuzzy string matching between each pair of answers $a_{1},a_{2}$ using Normalized Levenshtein Similarity (NLS):
@@
where $\text{editdistance}(a_{1},a_{2})$ computes the Levenshtein distance (Levenshtein et al., 1966) between the answers after tokenization. Two answers are considered equivalent if $\text{NLS}(a_{1},a_{2})\geq 0.75$.
Appendix C Extended Experimental Setup
@@ C.1 Dataset Descriptions
  • SlideVQA (Tanaka et al., 2023) is a multi-modal, multi-image VQA dataset consisting of $\geq$ 2600 slide decks that require single-hop, multi-hop, and numerical reasoning. Each slide deck corresponds to multiple questions. The license is available at  [^1] https://github.com/nttmdlab-nlp/SlideVQA/blob/main/LICENSE.
  • InfoVQA (Mathew et al., 2022) is a dataset that features questions that require reasoning and arithmetic skills. It contains over 30,000 questions with over 5,400 images. The dataset is released under CC-BY license.
@@ C.3 Implementation Details
We used EasyOCR [^2] https://github.com/JaidedAI/EasyOCR and Docling (Auer et al., 2024) for detecting textual and visual elements, respectively. Whenever applicable, the answers are generated using a temperature of 0.0 to ensure deterministic results.
Appendix D Additional Experiments
D.1 InfoVQA with Open-Source Backbones
Table 5 reports InfoVQA performance when SlideAgent is paired with open-source MLLMs, complementing the proprietary-model results in the main paper.
Model Overall Num F1
Raw LLMs
Llama-3.2-11B-Vision-Instruct 55.4 49.6 67.9
Phi-3-vision-128k-instruct 53.4 43.1 88.7
Qwen2.5-7B-Instruct 80.1 73.0 96.4
Qwen2.5-32B-Instruct 72.1 62.0 94.8
llava-1.5-7b-hf 25.1 10.6 83.9
llava-1.5-13b-hf 20.2 8.0 84.2
llava-v1.6-mistral-7b-hf 32.6 22.6 79.9
llava-v1.6-vicuna-13b-hf 34.9 24.8 80.9
InternVL3-8B 66.7 57.7 85.2
Multimodal RAG and Agentic Methods
ViDoRAG 67.1 59.2 85.9
SlideAgent 75.4 66.5 92.1
Impr. +8.7 +8.8 +6.8
Table 5: Performance comparison among open-source multimodal models and agentic methods.
D.2 Generalizability Across Models
SlideAgent’s framework is model-agnostic and provides consistent improvements when applied to stronger base models. Table 6 shows results with recent state-of-the-art models. With Gemini-2.5-Flash, SlideAgent achieves 90.5% overall accuracy (+6.7 points), 87.6% numeric (+9.3 points), and 93.6% F1 (+1.8 points) on SlideVQA, outperforming all models in Tables 2-4. With Qwen2.5-VL-7B (released Feb 2025), SlideAgent achieves 80.4% overall (+0.9 points), 73.8% numeric (+3.3 points), and 94.5% F1 (+0.2 points).
Model Overall Num F1
Gemini-2.5-Flash on SlideVQA
Raw Model 83.8 78.3 91.8
+ SlideAgent 90.5 87.6 93.6
Improvement +6.7 +9.3 +1.8
Qwen2.5-VL-7B-Instruct on SlideVQA
Raw Model 79.5 70.5 94.3
+ SlideAgent 80.4 73.8 94.5
Improvement +0.9 +3.3 +0.2
Table 6: SlideAgent with recent state-of-the-art models. Gemini-2.5-Flash + SlideAgent achieves best overall performance (90.5%), establishing state-of-the-art on SlideVQA.
D.3 Generalization to Other Document Types
To validate that SlideAgent’s hierarchical approach generalizes beyond slide decks, we evaluated on DocVQA (Mathew et al., 2021), a widely-used benchmark for document understanding with scanned documents, forms, and reports (Table 7). With GPT-4o as base model, SlideAgent achieves 94.5% overall accuracy (+0.9 points), 91.6% numeric (+2.5 points), and 97.5% F1 (+0.3 points) compared to the baseline. With InternVL3-8B, SlideAgent achieves 94.7% overall (+0.5 points), 93.4% numeric (+0.9 points), and 96.5% F1 (+0.5 points).
Model Overall Num F1
GPT-4o as base model
GPT-4o 93.6 89.1 97.2
SlideAgent 94.5 91.6 97.5
Improvement +0.9 +2.5 +0.3
InternVL3-8B as base model
InternVL3-8B 94.2 92.5 96.0
SlideAgent 94.7 93.4 96.5
Improvement +0.5 +0.9 +0.5
Table 7: Performance on DocVQA, demonstrating generalization beyond slide decks to scanned documents, forms, and reports.
D.4 Computational Cost and Scalability
An important aspect of SlideAgent’s design is that computational cost largely depends on the retriever used, which can be swapped based on latency requirements. Table 8 compares different retriever configurations versus single-shot GPT-4o.
Knowledge Construction (One-time, Amortized)
OCR/layout parsing takes $2.54\pm 0.32$ seconds per page (parallelized). Generating hierarchical knowledge ($\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}$) takes 51.5 seconds total for a typical slide deck, using $1544\pm 36$ MB memory. This is a one-time cost per document; the generated knowledge acts as a “cache” reusable across multiple queries.
Retrieval & QA (Per-Query)
[…]
Appendix E Use of AI Assistants
We used AI assistants to improve the paper’s presentation, including language polishing, clarity, and organization. AI assistants also provided support for data cleaning and processing. All technical experiments, analyses, and final content were verified by humans.
Model Latency (s) Tokens/sec Memory (MB) Input Tokens
Raw Model 9.12 $\pm$ 1.12 0.76 $\pm$ 0.27 1253 $\pm$ 35 15420 $\pm$ 2
SlideAgent (BM25) 9.44 $\pm$ 4.51 0.73 $\pm$ 0.61 1317 $\pm$ 32 1330 $\pm$ 212
SlideAgent (SFR) 19.17 $\pm$ 13.74 0.13 $\pm$ 0.05 2495 $\pm$ 7 1554 $\pm$ 342
Table 8: Computational cost comparison on SlideVQA. SlideAgent achieves >90% token reduction (15420 $\rightarrow$ 1330 tokens) while maintaining competitive latency. The one-time knowledge construction (51.5s total, 1544 MB) is amortized across multiple queries.
Appendix F Ethical Considerations
Data Privacy, Consent, and Intellectual Property.
Content Reliability and User Responsibility.
Figure 9: Performance comparison among variants of SlideAgent with base model InternVL3-8B.
@@ Appendix F Ethical Considerations
Notation Description
$p_{i},v_{i}$ A slide page and its visual content
$\mathcal{P}=\{p_{1},p_{2},\dots\}$ Set of pages in the slide deck
$q$ User query
$\hat{Q}$ Set of subqueries generated from the original query $q$
$a$ The final answer generated by SlideAgent
$b_{j}$ Bounding box coordinates of an element within a slide
$t_{j}$ Type of the element (e.g., text, chart, image, table)
[…]
Table 10: Mathematical notation used throughout the paper.
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Raw Models
Gemini 2.0 86.3 81.0 90.3 59.7 60.0 59.6 78.1 77.8 88.9
Gemini 2.5 89.0 85.7 93.2 61.5 65.0 59.5 76.6 76.4 83.3
Gemini 2.5-lite 81.8 75.5 90.1 56.3 55.0 57.0 78.1 78.4 66.7
Claude 4.1 85.7 82.4 89.7 58.0 77.5 48.3 52.6 52.0 60.8
Claude 3.5 58.2 64.0 50.9 55.8 83.7 42.5 47.8 48.0 40.3
GPT-4o 79.4 71.9 86.4 64.5 77.1 58.0 83.0 84.3 80.1
Multimodal RAG and Agentic Methods
ViDoRAG 81.8 73.8 87.0 65.8 78.0 58.7 84.2 84.7 81.5
SlideAgent 87.1 84.4 90.6 68.7 82.5 61.5 85.8 85.9 85.6
Impr. +7.7 +12.5 +4.2 +4.1 +5.4 +3.5 +2.8 +1.5 +5.5
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Raw Models
Llama-3.2-11B-Vision-Instruct 44.6 52.1 34.6 47.1 62.8 39.3 39.1 39.2 33.5
Phi-3-vision-128k-instruct 78.3 69.1 91.6 53.9 67.4 47.2 63.8 63.7 65.1
Qwen2.5-VL-7B-Instruct 85.1 77.7 95.3 57.7 69.8 51.8 52.7 52.0 77.8
Qwen2.5-VL-32B-Instruct 87.4 82.6 93.7 49.6 62.8 43.2 69.5 69.6 65.1
llava-1.5-7b-hf 42.9 27.9 79.9 24.6 15.0 39.4 14.2 14.4 20.9
llava-1.5-13b-hf 46.7 29.5 83.1 29.2 17.5 46.6 23.8 20.3 41.2
llava-v1.6-mistral-7b-hf 59.0 45.8 84.2 36.2 40.0 34.0 16.0 15.2 21.3
llava-v1.6-vicuna-13b-hf 62.3 48.2 87.1 54.7 62.5 49.5 35.2 30.7 67.6
InternVL3-8B 73.3 65.4 85.7 58.4 72.1 51.5 56.3 55.9 65.6
Baseline methods based on InternVL3-8B
ViDoRAG 76.3 68.1 89.9 61.3 75.1 53.9 58.4 58.7 67.3
SlideAgent 82.8 75.3 93.3 64.6 79.5 58.0 62.8 62.6 68.1
Impr. +9.5 +9.8 +7.6 +6.2 +7.4 +6.4 +6.5 +6.7 +2.5
Input Initial global knowledge $\mathcal{K}_{g}^{(0)}$ and concatenated page-level knowledge $\mathrm{concat}(\mathcal{K}_{p}^{1},\ldots,\mathcal{K}_{p}^{|\mathcal{P}|})$.
Instruction Rewrite the global knowledge from the full page-level evidence. Use $\mathcal{K}_{g}^{(0)}$ only as a weak prior for terminology and early context. Incorporate complete document coverage and narrative flow, including conclusions, summaries, and next steps that appear in later pages.
Update Rule Regenerate each field rather than appending snippets. If later-page evidence contradicts or refines an earlier hypothesis, overwrite the earlier field with the better-supported document-level interpretation.
Output Fields Title, objective, structure overview, key insights, audience, and tone.
Table 13: Prompt template for the global refinement operator. The operator performs a single-pass fieldwise rewrite from full-document page evidence.
### Cumulative Summary of the Preceding Slides {Cumulative Summary}
Element Type <<text | image | chart | table | icon | button | etc.>>
Position on Slide <<e.g., top-right, centered, below title>>
Semantic Role <<What is the element trying to do or communicate?>>
@@
Figure 10: Sample global knowledge $\mathcal{K}_{g}$ generated by the global agent, showing document-level summary, objectives, and narrative structure.
Figure 11: Sample page-level knowledge representation $\mathcal{K}_{p}^{i}$ generated by the page agent, capturing slide-specific content and cross-slide relationships.
Figure 12: Sample element-level knowledge representation $\mathcal{K}_{e}^{j}$ generated by the element agent.